# Observable Primordial Gravitational Waves from Non-minimally Coupled $R^{2}$ Palatini Modified Gravity

### v1

@@ Appendix D Matching the Mode Functions
$\beta_{+}$ $=-e^{\frac{i\pi}{4}(1+2\nu_{II})}\sqrt{\frac{\pi y_{\text{kin}}}{2}}~{}\alpha_{-}\left[-\left(H_{\nu_{II}-1}^{(1)}-H_{\nu_{II}+1}^{(1)}\right)k\eta_{\text{kin}}\right]$ (D.25)
$\beta_{+}=ie^{\frac{i\pi}{4}(1+2\nu_{II})}2^{2\nu_{II}}~{}\frac{H^{2}}{k^{2}}\frac{\Gamma(3/2-1/p)}{\Gamma(3/2)}(k\eta_{\text{kin}})^{1/p}$ (D.26)
@@
$v^{s}_{k}(\eta)=\frac{1}{\sqrt{2k}}\left[\gamma_{+}(k)e^{-ik\eta}+\gamma_{-}(k)e^{ik\eta}\right]$ (D.28)
@@
$y_{\text{reh}}\equiv kz_{\text{reh}}=k\left(\eta_{\text{reh}}-\frac{\eta_{\text{kin}}}{2}+\frac{1}{H}\right)\approx k\eta_{\text{reh}}$ (D.29)
Matching (LABEL:eq:modefn_match5) with (D.9) at $\eta_{\text{reh}}/y_{\text{reh}}$,
$\gamma_{+}e^{-ik\eta_{\text{reh}}}+\gamma_{-}e^{ik\eta_{\text{reh}}}=\sqrt{y_{\text{reh}}}\left[\beta_{-}rH_{\nu_{II}}^{(1)}(y_{\text{reh}})+\beta_{+}r^{\ast}H_{\nu_{II}}^{(2)}(y_{\text{reh}})\right]$ (D.30)
@@
$\begin{split}i\left(-\gamma_{+}e^{-ik\eta_{\text{reh}}}+\gamma_{-}e^{ik\eta_{\text{reh}}}\right)&=\frac{1}{2\sqrt{y_{\text{reh}}}}\left[\beta_{-}rH_{\nu_{II}}^{(1)}(y_{\text{reh}})+\beta_{+}r^{\ast}H_{\nu_{II}}^{(2)}(y_{\text{reh}})\right]\\ &+\sqrt{y_{\text{reh}}}\left[\beta_{-}r\frac{\mathrm{d}H_{\nu_{II}}^{(1)}}{\mathrm{d}y}(y_{\text{reh}})+\beta_{+}r^{\ast}\frac{\mathrm{d}H_{\nu_{II}}^{(2)}}{\mathrm{d}y}(y_{\text{reh}})\right]\end{split}$ (D.31)
@@
$\begin{split}\gamma_{-}&=\frac{e^{-ik\eta_{\text{reh}}}}{2}\left\{\beta_{-}r\left[H_{\nu_{II}}^{(1)}(y_{\text{reh}})\left(\sqrt{y_{\text{reh}}}-\frac{i}{2\sqrt{y_{\text{reh}}}}\right)-i\sqrt{y_{\text{reh}}}H_{\nu_{II}+1}^{(1)}(y_{\text{reh}})\right]\right.\\ &\left.+\beta_{+}r^{\ast}\left[H_{\nu_{II}}^{(2)}(y_{\text{reh}})\left(\sqrt{y_{\text{reh}}}-\frac{i}{2\sqrt{y_{\text{reh}}}}\right)-i\sqrt{y_{\text{reh}}}H_{\nu_{II}+1}^{(2)}(y_{\text{reh}})\right]\right\}\end{split}$ (D.32)
@@
$\gamma_{-}$ $=-\frac{i}{4\sqrt{y_{\text{reh}}}}\left(\beta_{-}rH_{\nu_{II}}^{(1)}(y_{\text{reh}})+\beta_{+}r^{\ast}H_{\nu_{II}}^{(2)}(y_{\text{reh}})\right)$ (D.33)
$=-\frac{i\beta_{-}}{2}e^{\frac{i\pi}{4}(1+2\nu_{II})}\frac{\Gamma(3/2)}{\Gamma(3/2-1/p)}\left(\frac{y_{\text{reh}}}{2}\right)^{\nu_{II}-1/2}$ (D.34)
$\gamma_{-}$ $=\frac{H^{2}}{k^{2}}\left(\frac{\eta_{\text{kin}}}{2\eta_{\text{reh}}}\right)^{1/p}$ (D.35)
@@
$\gamma_{+}=-\frac{H^{2}}{k^{2}}\left(\frac{\eta_{\text{kin}}}{2\eta_{\text{reh}}}\right)^{1/p}$ (D.36)

### v2

@@ Appendix E Matching the Mode Functions
$\beta_{+}$ $=-e^{\frac{i\pi}{4}(1+2\nu_{II})}\sqrt{\frac{\pi y_{\text{kin}}}{2}}~{}\alpha_{-}\left[-\left(H_{\nu_{II}-1}^{(1)}-H_{\nu_{II}+1}^{(1)}\right)k\eta_{\text{kin}}\right]$ (E.25)
$\beta_{+}=ie^{\frac{i\pi}{4}(1+2\nu_{II})}2^{2\nu_{II}}~{}\frac{H^{2}}{k^{2}}\frac{\Gamma(3/2-1/p)}{\Gamma(3/2)}(k\eta_{\text{kin}})^{1/p}$ (E.26)
@@
$v^{s}_{k}(\eta)=\frac{1}{\sqrt{2k}}\left[\gamma_{+}(k)e^{-ik\eta}+\gamma_{-}(k)e^{ik\eta}\right]$ (E.28)
@@
$y_{\text{reh}}\equiv kz_{\text{reh}}=k\left(\eta_{\text{reh}}-\frac{\eta_{\text{kin}}}{2}+\frac{1}{H}\right)\approx k\eta_{\text{reh}}$ (E.29)
Matching (LABEL:eq:modefn_match5) with (E.9) at $\eta_{\text{reh}}/y_{\text{reh}}$,
$\gamma_{+}e^{-ik\eta_{\text{reh}}}+\gamma_{-}e^{ik\eta_{\text{reh}}}=\sqrt{y_{\text{reh}}}\left[\beta_{-}rH_{\nu_{II}}^{(1)}(y_{\text{reh}})+\beta_{+}r^{\ast}H_{\nu_{II}}^{(2)}(y_{\text{reh}})\right]$ (E.30)
@@
$\begin{split}i\left(-\gamma_{+}e^{-ik\eta_{\text{reh}}}+\gamma_{-}e^{ik\eta_{\text{reh}}}\right)&=\frac{1}{2\sqrt{y_{\text{reh}}}}\left[\beta_{-}rH_{\nu_{II}}^{(1)}(y_{\text{reh}})+\beta_{+}r^{\ast}H_{\nu_{II}}^{(2)}(y_{\text{reh}})\right]\\ &+\sqrt{y_{\text{reh}}}\left[\beta_{-}r\frac{\mathrm{d}H_{\nu_{II}}^{(1)}}{\mathrm{d}y}(y_{\text{reh}})+\beta_{+}r^{\ast}\frac{\mathrm{d}H_{\nu_{II}}^{(2)}}{\mathrm{d}y}(y_{\text{reh}})\right]\end{split}$ (E.31)
@@
$\begin{split}\gamma_{-}&=\frac{e^{-ik\eta_{\text{reh}}}}{2}\left\{\beta_{-}r\left[H_{\nu_{II}}^{(1)}(y_{\text{reh}})\left(\sqrt{y_{\text{reh}}}-\frac{i}{2\sqrt{y_{\text{reh}}}}\right)-i\sqrt{y_{\text{reh}}}H_{\nu_{II}+1}^{(1)}(y_{\text{reh}})\right]\right.\\ &\left.+\beta_{+}r^{\ast}\left[H_{\nu_{II}}^{(2)}(y_{\text{reh}})\left(\sqrt{y_{\text{reh}}}-\frac{i}{2\sqrt{y_{\text{reh}}}}\right)-i\sqrt{y_{\text{reh}}}H_{\nu_{II}+1}^{(2)}(y_{\text{reh}})\right]\right\}\end{split}$ (E.32)
@@
$\gamma_{-}$ $=-\frac{i}{4\sqrt{y_{\text{reh}}}}\left(\beta_{-}rH_{\nu_{II}}^{(1)}(y_{\text{reh}})+\beta_{+}r^{\ast}H_{\nu_{II}}^{(2)}(y_{\text{reh}})\right)$ (E.33)
$=-\frac{i\beta_{-}}{2}e^{\frac{i\pi}{4}(1+2\nu_{II})}\frac{\Gamma(3/2)}{\Gamma(3/2-1/p)}\left(\frac{y_{\text{reh}}}{2}\right)^{\nu_{II}-1/2}$ (E.34)
$\gamma_{-}$ $=\frac{H^{2}}{k^{2}}\left(\frac{\eta_{\text{kin}}}{2\eta_{\text{reh}}}\right)^{1/p}$ (E.35)
@@
$\gamma_{+}=-\frac{H^{2}}{k^{2}}\left(\frac{\eta_{\text{kin}}}{2\eta_{\text{reh}}}\right)^{1/p}$ (E.36)